Solve the system by substition
x+y=12
y=2x

Answer by wallflower(4)

Hi there, for this problem you need to substitute y=2x for y in the x+y=12 and solve the equation.
x+2x=12
3x=12
x=4
y=2x so you have y=2(4) y=8
you can check your work by using x=4 and y=8 in the equations to see if they work.
4+8=12 so it works out.
Hope this helps you.